OpenStudy (anonymous):

what is the radius for x^2+(y-1)^2=12

OpenStudy (nenadmatematika):

square root of 12

OpenStudy (anonymous):

2V3

OpenStudy (anonymous):

this is the standard form: x^2+y^2=r^2 or (x+a)^ +(y+b)^2=r^2 such that a and b are constant that is why radius r is r^2=12 so r=12^.5= square root of 12